OpenStudy (anonymous):

what is n if √(3n+12)=√(n+8)?

OpenStudy (anonymous):

3n+12=n+8 2n=-4 n=-2

OpenStudy (ash2326):

square both sides 3n+12=n+8 3n-n=8-12 2n=-4 n=-2 let's check if both the root contains positive number 3n+12=6 n+8=6 so n=-2

OpenStudy (anonymous):

what is n if √n=√(2n-6)?

OpenStudy (anonymous):

n=2n-6 n=6

OpenStudy (anonymous):

what is x if √(11-x)=√(x-7)?

OpenStudy (anonymous):

11-x=x-7 18=2x x=9

OpenStudy (anonymous):

what is x if . √(72-x)=√(x/5)?

OpenStudy (anonymous):

72-x=x/5 360-5x=x 360=6x x=60

OpenStudy (anonymous):

what is x if √(x+3)=√(1-x)?

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

the same way. 1st hav to square both sides. then u get x+3=1-x 2x=-2 x=-1

OpenStudy (anonymous):

what is k if√(2k+40)=√(-16-2k)

OpenStudy (anonymous):

what is x √(x+8)=√(3x+8)?